OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences.

OpenStudy (anonymous):

A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

@ganeshie8 @TheSmartOne @pooja195 @poopsiedoodle @mathmate @uri @iambatman @inkyvoyd

ganeshie8 (ganeshie8):

Start by finding the starting price and growth rates of given stocks A, B, C

OpenStudy (anonymous):

sTOCK A 25

OpenStudy (anonymous):

STOCK B 22

OpenStudy (anonymous):

STOCK C 30

OpenStudy (anonymous):

idk dor stock d

ganeshie8 (ganeshie8):

we need to cook up stock D according to the given condition

OpenStudy (anonymous):

ok i think it is the second lowest

ganeshie8 (ganeshie8):

it must have the second lowest starting price, so how about making the starting price = `23` for stock D ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

what about growth rates of given stocks ?

OpenStudy (anonymous):

stock a 8 percent stock b idk stock c 4 percent and stock d 6 percent?

ganeshie8 (ganeshie8):

yes 6% looks good for stock D

OpenStudy (anonymous):

ok how about for stock b

ganeshie8 (ganeshie8):

you're ready to create the price function for stock D

ganeshie8 (ganeshie8):

looks like stick B is decreaesing so it must be having the lowest growth rate

OpenStudy (anonymous):

2 percent

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yes can you write a function for stock D ?

OpenStudy (anonymous):

yyes 6%

ganeshie8 (ganeshie8):

for stock D, does below work ?\[\large f(x) = 23(1.06)^x\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good, we are done with part 1

OpenStudy (anonymous):

ready to start second part @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@pooja195 @iambatman

OpenStudy (anonymous):

i am on part 2

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

the newprice function for stock D f(x) + g(x) will be 23(1.06)^x - 6

OpenStudy (cwrw238):

sorry can't stay gotta go right now

OpenStudy (anonymous):

@hartn @TuringTest @mathslover @mathmate

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@uri @bohotness

OpenStudy (anonymous):

@shrutipande9

OpenStudy (demonchild99):

sorry hun...i dont know how to do this @pooja195 can you help?

pooja195 (pooja195):

i wasnt sure about this either =/

OpenStudy (anonymous):

ok :(

OpenStudy (anonymous):

@ganeshie8

OpenStudy (demonchild99):

dont do that...then he wont help...cuz your spamming his name

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

@triciaal @Loser66

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Loser66

ganeshie8 (ganeshie8):

for part 2 use desmos to graph https://www.desmos.com/calculator/ut2tsxchvg

OpenStudy (anonymous):

idk how to do that

ganeshie8 (ganeshie8):

click that link

OpenStudy (anonymous):

ok is that the graph for part 2???

ganeshie8 (ganeshie8):

yes `Graph the price function for Stock D. Explain the key features of the graph with complete sentences.` see if you can identify any key features like y intercept etc.. and explain whatever you can

OpenStudy (anonymous):

umm may you help me w. that 2

OpenStudy (anonymous):

what is the y -intercept? and what does it represent

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

look at the graph https://www.desmos.com/calculator/lwe0xxrwan it seems the y intercept is (0, 23) ? looks it represents the starting value of stock D ?

OpenStudy (anonymous):

ok so that is what it represents the starting value for stock d? @ganeshie8

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

see if you can finish off the next part : `A recent drop in sales has affected Stock D with the function g(x) = -6. Explain to Gordon how Stock D s new price function, f(x) + g(x), will be created. Graph f(x) + g(x).`

OpenStudy (anonymous):

wait is that growth a function and what is the rate of the growth

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

6% right ?

OpenStudy (anonymous):

right but why is it a growth function

ganeshie8 (ganeshie8):

make a guess and tell me what you think

OpenStudy (anonymous):

because it is constantly increasing?

ganeshie8 (ganeshie8):

Exactly!

OpenStudy (anonymous):

may you help me w. the last part please @ganeshie8

ganeshie8 (ganeshie8):

`A recent drop in sales has affected Stock D with the function g(x) = -6. Explain to Gordon how Stock D s new price function, f(x) + g(x), will be created. Graph f(x) + g(x).` we have \(f(x) = 23(1.06)^x\) \(g(x) = -6\) \(f(x) + g(x) = ?\)

OpenStudy (anonymous):

fgx?

ganeshie8 (ganeshie8):

just add both the expressions

OpenStudy (anonymous):

i did

ganeshie8 (ganeshie8):

we have \(f(x) = 23(1.06)^x\) \(g(x) = -6\) \(f(x) + g(x) = 23(1.06)^x-6\) right ?

OpenStudy (anonymous):

right .

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

is this the correct graph https://www.desmos.com/calculator

ganeshie8 (ganeshie8):

https://www.desmos.com/calculator/ml8wtaxfeq

OpenStudy (anonymous):

thank you may you help me w. the last part please:)

ganeshie8 (ganeshie8):

``` Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(-x) f(x) + 3 ```

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

may u help me graph them

ganeshie8 (ganeshie8):

https://www.desmos.com/calculator/0zzwz6u4gk

OpenStudy (anonymous):

@ganeshie8 what is the transformation?

OpenStudy (anonymous):

@pooja195 may you help me explain the transformation

OpenStudy (anonymous):

@Loser66